OpenStudy (camerondoherty):

Check all polynomials that are perfect square trinomials: 9x2 − 6x + 1 36b2 − 24b + 8 16x2 + 24x + 9 4a2 − 10a + 25

OpenStudy (anonymous):

...

OpenStudy (anonymous):

look for ones that look like \[a^2+2ab+b^2=(a+b)^2\] for example \[9x^2-6x+1=(3x-1)^2\]

OpenStudy (camerondoherty):

does it need to have the + and - in order?

OpenStudy (anonymous):

also \[ 16x^2 + 24x + 9 =(5x+3)^2\]

OpenStudy (anonymous):

no

OpenStudy (camerondoherty):

so they all work?

OpenStudy (anonymous):

\[36b^2 − 24b + 8 \] is not one because \(8\) is not a perfect square

OpenStudy (anonymous):

I would choose 9x^2 − 6x + 1 right away, knowing that it is the same as x^2-6x+9, but reversed, if you know what I mean.

OpenStudy (camerondoherty):

oh ok thank you!

OpenStudy (anonymous):

neither is \(4a^2 − 10a + 25 \)

OpenStudy (camerondoherty):

y?

OpenStudy (anonymous):

it looks like one, but \((2a-5)^2=4a^2-20a+25\)

OpenStudy (anonymous):

if \(a=2x,b=-5\) then \(2ab=-20x\) not \(-10x\) it was a trick

OpenStudy (camerondoherty):

oh ok lol

OpenStudy (camerondoherty):

Thank you!

OpenStudy (anonymous):

yw